A stock has a beta of 1.5 and is expected to pay a $2 dividend per share next year. The dividend is expected to grow at 1% per year. The risk-free rate is 1%, and market risk premium is 6%. What should be the fair stock price?
Answer #1

Required return=risk free rate+beta*market risk premium

=1+(1.5*6) =10%

Hence stock price=D1/(required return-growth rate)

=2/(0.1-0.01)

=$22.22(approx).